OpenStudy (lovelyharmonics):

A solution is made by dissolving 25.5 grams of glucose (C6H12O6) in 398 grams of water. What is the freezing-point depression of the solvent if the freezing point constant is -1.86 °C/m? Show all of the work needed to solve this problem.

OpenStudy (aaronq):

you would use \(\Delta T=i*m*K_f\) where m is molality. If you want i can expand this formula \(\Delta T=\dfrac{m*K_f}{M*kg_{solvent}}\) m= mass, M=molar mass

OpenStudy (lovelyharmonics):

Wait so m=25.5 how do I get molar mass?

OpenStudy (aaronq):

You would add up the individual molar masses of C, O and H in the right proportions (6 times carbon, 6 times oxygen and 12 times hydrogen).

OpenStudy (lovelyharmonics):

Okay so carbon would be 12*6=72, oxygen would be 16*6=96, and hydrogen would be 12*1=12 so it would total 180 2.24*K_f / 180*kg_solvent ?

OpenStudy (aaronq):

you've got the mass wrong \(\Delta T=\dfrac{(25.5~g)*K_f}{ (180~g/mol)*kg_{solvent}} \)

OpenStudy (lovelyharmonics):

Oh .-.i didn't mean for it to be a 4. Wait, what stands for K_f and kg?

OpenStudy (aaronq):

Kf is the freezing point depression constant, in this case it's -1.86 °C/m Kg is the kilograms of solvent (in this case water)

OpenStudy (lovelyharmonics):

Okay so it's 25.5*-1.86 / 180*398 23.64/71640

OpenStudy (aaronq):

you didn't convert the mass of water to kilograms

OpenStudy (lovelyharmonics):

Oh okay so that would be 0.398*180=71.64 So 23.64/71.64

OpenStudy (lovelyharmonics):

That's a really weird number.... So I guess it's 0.33 @aaronq

OpenStudy (aaronq):

i got \(\Delta T\)=-0.6620603015075377 =-0.66

OpenStudy (aaronq):

Because the freezing point of water is 0 Celsius, your answer is -0.7 Celsius

OpenStudy (lovelyharmonics):

Wait did I calculate it wrong? Because 23.64/71.64=0.33

OpenStudy (aaronq):

you got the numerator wrong, it's 25.5*-1.86=-47.43

OpenStudy (aaronq):

you subtracted, you were supposed to multiply it

OpenStudy (lovelyharmonics):

Thank you c:

OpenStudy (aaronq):

no problem !